# Large language model validity via enhanced conformal prediction methods

### v1

@@ B.1 Method for learning $\alpha(\cdot)$
As discussed in the main text, when running our method we split our training data into a portion for fitting $\alpha(\cdot)$ and portion for calibrating the conformity score. Then, we further divide the portion of the data that is used to fit $\alpha(\cdot)$ into two folds. We use the first fold to run the conditional conformal procedure at a fixed grid of quantiles, $\alpha\in\{0.01,\dots,0.99\}$, and the second dataset to evaluate $Q(\mathbf{C}_{i};\hat{\tau}_{\alpha}(X_{i}))$ for each choice of $\alpha$. Here, $\hat{\tau}_{\alpha}(X_{i})$ denotes the conditional conformal cutoff evaluated for test claims $\mathbf{C}_{i}$ at fixed level $1-\alpha$.
After collecting our estimates of $Q(\mathbf{C}_{n+1};\hat{\tau}_{\alpha}(X_{n+1}))$, we compute the maximum $\alpha$ at which the quality criterion is achieved. Formally, for each point $(\mathbf{C}_{i},X_{i})$ in the dataset held out from fitting, we compute
@@ Appendix C Details of the Boosting Procedure
For the boosting procedure, we will assume that we are working with the derandomized variant of the conditional conformal method of Gibbs et al. [10]. In addition, these results will hold only for the finite-dimensional function class variant of their method, i.e., we assume that
=$\mathcal{F}=\{\Phi(\cdot)^{\top}\beta:\beta\in\mathbb{R}^{d}\}.$
@@
Assume that there the solution to the quantile regression given by (7) with $S=\hat{\tau}_{i}(\theta)$ is non-degenerate, and let $B$ denote the optimal basis of this solution. Then $\hat{\tau}_{i}(\theta)$ is locally linear in $\theta$ with derivative given by
$\partial_{\theta}\hat{\tau}_{i}(\theta)=\Phi(X)_{B}^{-1}\partial_{\theta}S_{B}(\theta).$ (14)
Proof    If the primal solution is non-degenerate, it is well-known that the dual variables corresponding to the optimal basis are strictly bounded away from its constraints [8]. Since the dual variables are piecewise linear in $\{S_{i}\}_{i=1}^{n}$ (a fact commonly exploited in sensitivity analysis for linear programs) [8], we know that there exists some $\epsilon$-ball in $S$ for which the basis is unchanged. Therefore, so long as $S$ is continuous in $\theta$ (and we have assumed differentiability), the stated result holds. ∎
When running Algorithm 1, we must set aside some data to run the final conformal procedure. Once the scores are boosted, they must be evaluated on a fresh dataset for the resulting filtered output to satisfy any of our guarantees.
Data: Boosting dataset $\mathcal{D}=\{\mathbf{D}_{i}\}_{i=1}^{n}$, conformity score function $S_{\theta}(\cdot,\cdot)$, function class basis $\mathbf{\Phi}$, initialization $\theta_{0}$, sigmoid temperature $\lambda$, level $\alpha$, boosting iterations $N$.
for $i\in[N]$ do
@@
       $\mathcal{D}_{1},\mathcal{D}_{2}=\textsf{RandomSplit}(\mathcal{D})$ ;
       $\Phi=\Phi(\mathcal{D}_{1})$ ;
        /* Shorthand for computing the basis on the first set */
       $S=S_{\theta}(\mathcal{D}_{1})$;
        /* Shorthand for computing the scores on the first set */
       for $\mathbf{D}_{i}\in\mathcal{D}_{2}$ do
             conf = CondConf$(S_{\theta},\Phi)$;
             $B=$ conf.getOptimalBasis$(\alpha,\mathbf{D}_{i})$;
             $\hat{\beta}_{i}=\Phi_{B}^{-1}(\mathcal{D}_{1})S_{B}$;
             $\ell=\ell+\sigma_{\lambda}(\Phi_{n+1}^{\top}\hat{\beta}_{i}-\sum_{j=1}^{n_{i}}p_{\theta}(P_{i},C_{ij}))$;
      $\theta_{i}=\text{Adam}(\ell)$.step();
return $\theta_{N}$.
@@
We make one important simplification in this algorithm to produce this paper’s experiments. The subroutine conf.getOptimalBasis obtains the optimal basis in Algorithm 1 by identifying the $\text{dim}(\Phi)$ points that are interpolated by the regression when $S_{n+1}>\hat{\tau}_{i}(\theta)$. Identifying this set is computationally burdensome since each gradient step in $\theta$ requires computing $\hat{\tau}_{i}(\theta)$ (and therefore $\hat{\beta}_{i}$) for each point in the second split. Thus, the gradient computation step scales quadratically in $|\mathcal{D}_{2}|$. A linear-time approximation to the gradient can be obtained by defining $B$ as the optimal basis from the quantile regression fit to $\mathcal{D}_{1}$ alone. This approximation, which does not account for the test-point augmentation in the conditional conformal procedure, is much faster since it yields a single $\hat{\beta}$ to use for all the test points. To concretely modify Algorithm 1, $B$ is defined as the optimal basis of the quantile regression fit to $\mathcal{D}_{1}$ alone, the computation in the nested for loop is removed, and the estimated threshold in the sigmoid function becomes $\Phi_{n+1}^{\top}\hat{\beta}$.
Appendix D Experiments
D.1 Claim scores
@@
We collect several features to use for estimating a confidence level in each claim. Unless otherwise noted, the claim scoring function used in this paper is defined by the “frequency” score examined by Mohri and Hashimoto [19]. We compute this score by querying GPT-3.5-Turbo 5 times at a temperature of $T=5$. For each claim, we re-query the LLM to ask if the claim is supported in the resampled generation. These scores which can be $-1$ for contradicted, 00 for not present, and $1$ for supported are then averaged over the $5$ responses. For example, a claim that is also present in $3$ out of the $5$ responses, but absent in the other two would receive a score of $0.6$.
@@
Since the frequency score requires many API queries, we also investigate several cheaper and faster-to-compute features for claim scoring. Following Mohri and Hashimoto [19], we also collect the self-reported probability obtained by directly asking the LLM to report its estimated probability of claim correctness to three significant figures. We also compute the ordinal score, i.e., the order in which the claim appeared in the original response. Going beyond Mohri and Hashimoto [19], we compute the LLM’s internal probability of each claim by querying the model with the list of claims and asking it to output a single-character (T or F) assessment of its beliefs. The internal probability of the model can then be computed by exponentiating the log probability of the returned token.
D.2 Prompts
@@
To parse the initial response into scorable sub-claims, we prompt the LLM (GPT-4o, in our experiments). After first parsing the prompt into its component sentences, we copy the method of Min et al. [18] (made available on GitHub under an MIT license) and use in-context learning to improve the parsing accuracy of the LLM. We first use the BM25 ranking function to match our sentences to similar examples previously parsed by Min et al. [18]. We prepend these gold-standard examples as well as the new sentence to be parsed to the prompt,
@@
D.3 Details of Section 4.3
Here we also provide some additional details regarding the function classes used in our biography experiments. In all examples, we aim to provide uniform coverage guarantees regardless of the underlying popularity of the topic. In the fixed-level filtered biographies, we run the conditional conformal method with =$\mathcal{F}=\{\beta_{0}+\sum_{i=1}^{3}\beta_{i}v^{i}\mid\beta\in\mathbb{R}^{4}\}$, where $v$ denotes the number of views of the corresponding Wikipedia page in January 2023. In the level-adaptive examples, a nearly identical two-step workflow is used to estimate $\alpha(\cdot)$ and estimate the method’s coverage. Here we make one small change; we estimate $\alpha(\cdot)$ using a $0.25$-quantile regression over $\mathcal{F}$ (see Figure 7 for a plot of this estimate). In the final level-adaptive fit, we add an additional linear term to this function class given by $\beta_{4}(\alpha(\cdot)-0.1)^{2}$. Note that in our experiments, we truncate $\alpha(\cdot)$ to lie above $0.1$. As a result, a substantial proportion of outputs receive that estimated quantile and thus, $(\alpha(\cdot)-0.1)^{2}$ correlates with the more variable portion of the quantile function.
Though the final experiments in this section are run using the more performant frequency score, the right panel of Figure 5 is obtained by finding an optimal linear combination of the three cheaper-to-obtain claim scores (self evaluation, token probability, ordinal). We obtain the displayed result by running $1000$ steps of Adam (with PyTorch defaults, i.e., learning rate $0.001$) through the conditional conformal procedure given by the same $\mathcal{F}$ described in the previous paragraph.

### v2

@@ B.1 Method for learning $\alpha(\cdot)$
As discussed in the main text, when running our method we split our training data into a portion for fitting $\alpha(\cdot)$ and portion for calibrating the conformity score. Then, we further divide the portion of the data that is used to fit $\alpha(\cdot)$ into two folds. We use the first fold to run the conditional conformal procedure at a fixed grid of quantiles, $\alpha\in\{0.01,\dots,0.99\}$, and the second dataset to evaluate $Q(\mathbf{C}_{i};\hat{\tau}_{\alpha}(X_{i}))$ for each choice of $\alpha$. Here, $\hat{\tau}_{\alpha}(X_{i})$ denotes the conditional conformal cutoff evaluated for test claim $\mathbf{C}_{i}$ at fixed level $1-\alpha$.
After collecting these data, we compute the maximum $\alpha$ at which the quality criterion is achieved. Formally, for each point $(\mathbf{C}_{i},X_{i})$ in the dataset held out from fitting, we compute
@@ Appendix C Details of the Boosting Procedure
For the boosting procedure, we will assume that we are working with the derandomized variant of the conditional conformal method of Gibbs et al. [10]. In addition, these results will hold only for the finite-dimensional function class variant of their method with no regression penalty, i.e., in the notation of Appendix A we assume that
=$\mathcal{F}=\{\Phi(\cdot)^{\top}\beta:\beta\in\mathbb{R}^{d}\},\qquad\mathcal{P}(\cdot)=0.$
@@
Let $\hat{\tau}_{i}(\theta)$ denote the non-randomized filtering threshold defined via (8). Assume that the threshold is finite and that the augmented quantile regression for all $S>\hat{\tau}_{i}(\theta)$ admits a unique non-degenerate basic solution. Let $B$ denote this basis. Then, $\hat{\tau}_{i}(\theta)$ has derivative
$\partial_{\theta}\hat{\tau}_{i}(\theta)=\Phi(X_{n+i})^{\top}\left(\Phi(X)_{B}^{-1}\partial_{\theta}S_{B}(\theta)\right).$ (14)
Proof    Assume without loss of generality that $i=1$. We drop the subscript on $\hat{\tau}$ for notational convenience. To differentiate between the imputed value of the $(n+1)$-st score and the vector of all observed scores, we will use $S$ to refer to the former and $\mathbf{S}$ to the latter. When the optimal basis is unique, we will assume that the primal solution is defined via the interpolator of the basis. Otherwise, we assume it is obtained via some symmetric algorithm.
First, we establish that $\theta\mapsto\Phi(X_{n+1})^{\top}\beta_{S}(\theta)$ is locally linear in $\theta$ for $S>\hat{\tau}(\theta)$. First, observe that the non-interpolated points of the quantile regression correspond to the non-basic variables in the solution, and that their residuals are equivalent to the non-basic “reduced costs” of the LP. Standard results from LP sensitivity analysis establish that when these reduced costs are bounded away from 00 (implied by our assumptions of non-degeneracy and uniqueness), the basis of the LP is unchanged under sufficiently small perturbations of the objective coefficients (i.e., for perturbations of $\theta$) [8]. Since $\beta_{S}$ is defined by the solution to a linear system of equations involving the basis, the local constancy of the basis implies the claim of local linearity. Second, we establish that $S\mapsto\beta_{S}(\theta)$ for $S>\hat{\tau}(\theta)$ is constant. This is immediate from the KKT conditions for (7). The optimal basis is unchanged for all $S>\hat{\tau}(\theta)$.
Fix any $\theta_{0}$. We claim that there exists a ball around $\theta_{0}$ such that $\hat{\tau}(\theta)=\Phi(X_{n+1})^{\top}\beta_{\hat{\tau}(\theta_{0})+1}(\theta)$.
To prove this, first note that we know that for $\theta$ sufficiently close to $\theta_{0}$, $\hat{\tau}(\theta_{0})+1>\Phi(X_{n+1})^{\top}\hat{\beta}_{\hat{\tau}(\theta_{0})+1}(\theta)$. Since $S\mapsto\beta_{S}(\theta)$ is constant above $\hat{\tau}(\theta)$, this immediately implies that for $S>\Phi(X_{n+1})^{\top}\hat{\beta}_{\hat{\tau}(\theta_{0})+1}(\theta)$, $\hat{\beta}_{S}(\theta)=\hat{\beta}_{\hat{\tau}(\theta_{0})+1}(\theta)$ and thus $S>\Phi(X_{n+1})^{\top}\hat{\beta}_{S}(\theta)$. Hence, $\hat{\tau}(\theta)\leq\Phi(X_{n+1})^{\top}\hat{\beta}_{\tau(\theta_{0})+1}(\theta)$.
Now, assume by contradiction that $\hat{\tau}(\theta)<\Phi(X_{n+1})^{\top}\hat{\beta}_{\tau(\theta_{0})+1}(\theta)$. For ease of notation let $v=\Phi(X_{n+1})^{\top}\hat{\beta}_{\tau(\theta_{0})+1}(\theta)$. Now, since $\hat{\tau}(\theta)<v$ we must have $v>\Phi(X_{n+1})^{\top}\hat{\beta}_{v}(\theta)$. Recall that we additionally know that $\hat{\tau}(\theta_{0})+1>\Phi(X_{n+1})^{\top}\hat{\beta}_{\hat{\tau}(\theta_{0})+1}(\theta)$. So, by our assumption that $S\mapsto\beta_{S}(\theta)$ is constant above $\hat{\tau}(\theta)$ we find that $\hat{\beta}_{\tau(\theta_{0})+1}(\theta)=\hat{\beta}_{v}(\theta)$. Thus, $v=\Phi(X_{n+1})^{\top}\hat{\beta}_{\tau(\theta_{0})+1}(\theta)=\Phi(X_{n+1})^{\top}\hat{\beta}_{v}(\theta)$, which gives the desired contradiction.
The above proves that $\tau(\theta)=\Phi(X_{n+1})^{\top}\beta_{\tau(\theta_{0})+1}(\theta)$ locally in a ball around $\theta_{0}$. The desired result follows by our previous proof of the differentiability of the right hand-side.
With this result in hand, our boosting procedure is now given in detail in Algorithm 1 below. In this algorithm we use the notation $\Phi(I)=[\Phi(X_{i})]_{i\in I}$ and $S_{\theta}(I)=(S_{\theta}(\mathbf{C}_{i},\mathbf{W}_{i}))_{i\in I}$ to refer the feature matrix and scores for the data points in $I$.
Data: Boosting dataset $\mathcal{D}_{\text{boost}}=\{\mathbf{D}_{i}\}_{i=1}^{m}$, conformity score function $S_{\theta}(\cdot,\cdot)$, function class basis $\mathbf{\Phi}$, initialization $\theta_{0}$, sigmoid temperature $\lambda$, level $\alpha$, boosting iteration count $T$.
for $t\in[T]$ do
@@
       $\mathcal{D}_{1},\mathcal{D}_{2}=\textsf{RandomSplit}(\mathcal{D}_{\text{boost}})$ ;
       for $i\in\mathcal{D}_{2}$ do
             $B=$ getOptimalBasis$(\alpha,\Phi(\mathcal{D}_{1}),S_{\theta}(\mathcal{D}_{1}),\Phi(\{i\}),S_{\theta_{t-1}}(\{i\}))$;
             $\hat{\beta}_{i}=\Phi_{B}^{-1}(\mathcal{D}_{1}\cup\{i\})S_{B}(\mathcal{D}_{1}\cup\{i\})$;
             $\ell=\ell+\sum_{j=1}^{k_{i}}\sigma_{\lambda}(\Phi(\{i\})^{\top}\hat{\beta}_{i}-p_{\theta}(P_{i},C_{ij}))$;
      $\theta_{t}=\text{Adam}(\ell,\theta_{t-1})$.step();
return $\theta_{T}$.
@@
In Algorithm 1 above, $\textsf{getOptimalBasis}(\cdot)$ refers to the subroutine that computes the optimal basis $B$ appearing in the statement of Proposition C.1 for text point $i$.
@@
In order to decrease computational complexity, the experiments in this paper are run using a simplified version of Algorithm 1. In particular, the subroutine getOptimalBasis obtains the optimal basis in Algorithm 1 by identifying the $\text{dim}(\Phi)$ points that are interpolated by the regression when $S>\hat{\tau}_{i}(\theta)$. Identifying this set is computationally burdensome since each gradient step in $\theta$ requires computing $\hat{\tau}_{i}(\theta)$ (and therefore $\hat{\beta}_{i}$) for each point in the second split. Thus, the gradient computation step scales quadratically in $|\mathcal{D}_{2}|$. A linear-time approximation to the gradient can be obtained by defining $B$ as the optimal basis from the quantile regression fit to $\mathcal{D}_{1}$ alone. This approximation, which does not account for the test-point augmentation in the conditional conformal procedure, is much faster since it yields a single $\hat{\beta}$ to use for all the test points. The experiments in this paper are run using this simplified procedure. To concretely modify Algorithm 1, $B$ is defined as the optimal basis of the quantile regression fit to $\mathcal{D}_{1}$ alone, the computation in the nested for loop is removed, and the estimated threshold in the sigmoid function becomes $\Phi^{\top}(\{i\})\hat{\beta}$.
Appendix D Additional experiments
D.1 Additional results for Section 4.2
Additional details for the experiment set-up
The function class $\mathcal{F}$ is defined by the linear combination of an intercept, the number of characters in the prompt, the number of characters in the response, the mean frequency score assigned to the claims, the standard deviation of the frequency scores assigned to the claims, and group-indicators corresponding to the source dataset. The results shown in Figure 2 are obtained by running the conditional boosting algorithm for $1000$ steps using the Adam optimizer with learning rate set to $0.001$.
We estimate $\alpha(\cdot)$ by dividing the first split of the data into two further sub-parts. On the first half of this further subdivisionx ($n=720$), we run the fixed-level conditional conformal procedure (with the same $\mathcal{F}$ used to boost the scores) for each $\alpha$ in a grid of $50$ evenly spaced values between 00 and $1$. Using the second half of this dataset, we evaluate the percentage of retained claims at each $\alpha$. We record the smallest $\alpha$ for which at least $70\%$ of claims are preserved at all larger values of $\alpha$. We then obtain our estimate of $\alpha(\cdot)$ by fitting a $0.85$-quantile regression over $\mathcal{F}$ to this dataset. For interpretability and numerical stability in the next step, the final output of the function is truncated to lie between $0.1$ and $0.5$.
Additional experiments
Figure 9 decomposes the effect of each method in this paper on claim retention and calibration for the MedLFQA benchmark. For a fixed-level guarantee, we observe that the boosting method substantially improves claim retention. When we adapt the level to guarantee high claim retention, the improvement of the boosting method is visible in the second panel of Figure 9. Namely, we find that when the level-adaptive method is augmented with boosted scores we obtain the same claim retention while issuing stronger guarantees as compared to the un-boosted scores. Finally, the third panel of Figure 9 verifies that our reported values of $1-\alpha(X_{n+1})$ (approximately) match the realized error rates.
Figure 9: The left panel of the figure shows the percentage of claims retained by various methods on the MedLFQA benchmark of Jeong et al. (2024). This benchmark contains many datasets each of which is displayed on the x-axis. Before boosting, we define our claim score by an equally-weighted ensemble of previously published claim scoring functions. We run $100$ trials in which we resample a boosting/$\alpha(\cdot)$-estimation split ($n=1441$), calibration split ($n=2354$), and test split ($n=500$). We plot $4$ filtering methods: the fixed-level ($0.91-\alpha=$ 0.9) method that guarantees conditional validity over the function class given by dataset indicators and prompt metadata (prompt length, response length, as well as the mean and standard deviation of the “log probability” claim score over the output) (blue), the same conditionally valid method with adaptive level $\alpha(X_{n+1})$ (orange), the conditionally valid method with boosted scores (green), and a combination method that incorporates both conditional-boosting and level-adaptive CP (red). All methods are set-up to ensure that the final output contains $\mathbf{0}$ false claims. The middle panel shows that boosting allows the level-adaptive procedure to issue guarantees with higher confidence. The right panel verifies that the reported nominal levels $1-\alpha(X_{n+1})$ match the empirical error frequencies over equi-spaced bins of width $0.05$; these bin indicators are included in the function class used in the level-adaptive procedure.
Using the same experimental set-up as the ablations, Figure 10 compares the effect of the choice of claim scoring method on the nominal guarantee offered by the level-adaptive procedure. Note that the log-probability and frequency claim scores appear to offer the strongest set of guarantees. However, the markedly lower level of claim retention for the log-probability score make the guarantees difficult to directly compare.
Figure 10: The set-up is identical to that of Figure 9. Here we do not consider boosted scores, but display the performance of the level-adaptive method for each claim scoring approach in isolation.
D.2 Additional results for Section 4.3
Additional details for the experiment set-up
Here, we also provide some additional details regarding the function classes used in our biography experiments. In all examples, we aim to provide uniform coverage guarantees regardless of the underlying popularity of the topic. In the fixed-level filtered biographies, we run the conditional conformal method with =$\mathcal{F}=\{\beta_{0}+\sum_{i=1}^{3}\beta_{i}v^{i}\mid\beta\in\mathbb{R}^{4}\}$, where $v$ denotes the number of views of the corresponding Wikipedia page in January 2023. In the level-adaptive examples, we use a nearly identical workflow to the previous section. The only difference is that here we estimate $\alpha(\cdot)$ using a $0.75$-quantile regression over $\mathcal{F}$ (see Figure 7 for a plot of this estimate). Additionally, in the final level-adaptive fit, we add a linear term to this function class given by $\beta_{4}(\alpha(\cdot)-0.1)^{2}$. Note that in our experiments, we truncate $\alpha(\cdot)$ to lie above $0.1$. As a result, a substantial proportion of outputs receive that estimated quantile and thus, $(\alpha(\cdot)-0.1)^{2}$ correlates with the more variable portion of the quantile function.
Though the first two panels of Figure 5 and Figure 6 are run using the more performant frequency score, the right panel of Figure 5 is obtained by finding an optimal linear combination of the three cheaper-to-obtain claim scores (self evaluation, token probability, ordinal). We obtain the displayed result by running $1000$ steps of Adam (with PyTorch defaults, i.e., learning rate $0.001$) through the conditional conformal procedure given by the same $\mathcal{F}$ described in the previous paragraph.
Additional experiments
Figures 11 and 12 decompose the effect of each method on claim retention and calibration for the FActscore benchmark. Notably, we observe that optimally ensembling the 3 cheaper-to-compute claim scores using conditional boosting improves claim retention and/or shifts the distribution of issued guarantees to the right. Furthermore, regardless of the guarantee issued, the third panel of both of these figures confirms that our method is well-calibrated. Finally, as the second panel of Figure 12 confirms, if we aim to guarantee 00 false claims in the final output, it is unlikely that we issue a nominal probability above $50\%$. This motivates our choice to instead guarantee that the filtered output contains at most three false claims.
Figure 11: This figure replicates the previous ablation study (namely Figure 9) for the FActscore dataset of Min et al. [20]. For the sake of brevity, we only describe the differences compared to Figure 9. The frequency of each biography topic varies; the displayed groups correspond to Wikipedia view counts in Jan. 2023 that are binned into the intervals $[1000000,\infty),[100000,1000000),[1000,100000),[100,1000),[0,100)$. The function class used for calibration is defined by the linear combination of these bin indicators and the view count of each Wikipedia article. We run $100$ trials in which we resample a boosting/$\alpha(\cdot)$-estimation split ($n=847$), calibration split ($n=5338$), and test split ($n=500$). All methods are set-up to ensure that (with high probability) the final output contains no more than $3$ false claims.
Figure 12: The set-up is identical to that of Figure 11, but now all methods are set-up to ensure that (with high probability) the final output contains no more than 00 false claims.
Using the same experimental set-up as the ablations, Figures 14 and 13 compare the effect of the choice of claim scoring method on the nominal guarantee offered by the level-adaptive procedure. Note that once again the log-probability and frequency claim scores appear to offer the strongest set of guarantees.
Figure 13: The set-up is identical to that of Figure 11. Here we do not consider boosted scores, but display the performance of the level-adaptive method for each claim scoring approach in isolation.
Figure 14: The set-up is identical to that of Figure 12. Here we do not consider boosted scores, but display the performance of the level-adaptive method for each claim scoring approach in isolation.
Appendix E Additional scoring details
E.1 Claim scores
@@
We collect several features to use for estimating a confidence level in each claim. Unless otherwise noted, the claim scoring function used in this paper is defined by the “frequency” score examined by Mohri and Hashimoto [21]. We compute this score by querying GPT-3.5-Turbo 5 times at a temperature of $T=5$. For each claim, we re-query the LLM to ask if the claim is supported in the resampled generation. These scores which can be $-1$ for contradicted, 00 for not present, and $1$ for supported are then averaged over the $5$ responses. For example, a claim that is present in three out of the five responses, but absent in the other two would receive a score of $0.6$.
@@
Since the frequency score requires many API queries, we also investigate several cheaper and faster-to-compute methods for claim scoring. Following Mohri and Hashimoto [21], we also collect the self-reported probability obtained by directly asking the LLM to report its estimated probability of claim correctness to three significant figures. We also compute the ordinal score, i.e., the order in which the claim appeared in the original response. Going beyond Mohri and Hashimoto [21], we compute the LLM’s internal probability of each claim by querying the model with the list of claims and asking it to output a single-character (T or F) assessment of its beliefs. The internal probability of the model can then be computed by exponentiating the log probability of the returned token.
E.2 Prompts
@@
To parse the initial response into scorable sub-claims, we prompt the LLM (GPT-4o, in our experiments). After first parsing the initial response into its component sentences, we copy the method of Min et al. [20] (made available on GitHub under an MIT license) and use in-context learning to improve the parsing accuracy of the LLM. We first use the BM25 ranking function to match our sentences to similar examples previously parsed by Min et al. [20]. We prepend these gold-standard examples as well as the new sentence to be parsed to the prompt,